1.Your profit P is determined by subtracting the cost C (the amount of money it costs to operate a business) from the revenue R (the amount of money you earn from selling your product). Profit can be represented algebraically by the equations:
Profit=Revenue-Cost
OR
P = R - C
a)Rewrite the formula to solve for C.
b)Suppose your profit for one day is $1,281, and the cost of running the business for the day is $1,587. What is the revenue for that day? Explain your answer.

Answer by jorel1380(3719)

P=R-C
--------------
a)C=R-P
----------
b)1281=R-1587
R=1281+1587=$2868..